What is the balanced chemical equation of 55ml of 0.150M Copper (II)Sulfate and 48ml of .250M Sodium Hydroxide?

What is the limiting and excess reactants?
1
Expert's answer
2019-01-22T13:35:53-0500

Balanced chemical equation:

CuSO4(aq) + 2NaOH(aq) = Cu(OH)2(s) + Na2SO4

n = c*V

n(CuSO4) = 0.150 mol/L *0.055 L = 0.00825 mol

n(NaOH) = 0.250 mol/L * 0.048 L = 0.012 mol

According to equation mole ratio n(CuSO4): n(NaOH) = 1:2. Amount os substance of NaOH should be twice the amount of CuSO4.

Amount of NaoH should be: n(NaOH) = 2*n(CuSO4) = 2*0.00825 = 0.0165 mol.

But we have n(NaOH) = 0.012 mol.

0.012 mol < 0.0165 mol, then NaOH is the limiting reactant. CuSO4 is the excess reactant.